CH/Tech/92/07

Update

Silicon Detectors

Author: Brian Fulton

Date: 19 May 1992

Distribution: WNC,NMC,PVD,BRF,GJG,RAH,JSL,VP,WDMR,GT,DLW

Silicon Detector Developments

Developments since the last meeting are summarised below

• Q-Par Angus First prototype detector (1 kΩ □^{−1}) delivered and ready for tests. The second prototype (2 kΩ □^{−1}) due in two weeks time.

• Micron Double sided strip detector (NASA flight test module) awaiting tests. Colin Wilburn has been in touch to offer more details on this (sending a copy of a report presented at a conference at Santa Fe).

• Canberra Quotation received from Canberra on their PAD detectors (our PSSSD). Specification looks similar to Micron but cost is higher. Michel Ceuppens, one of their sales and marketing people from Belgium, called in for a discussion a few weeks ago as he was in the country. He took back a number of technical enquiries for Paul Berger. Paul Berger will be here for a meeting on 10 June to discuss these points. It seems they are very interested in working towards either design (the sheet resistance readout or double sided design with inline resistors). Had a chat with Siegfried Hoffman at GSI who uses some of their detectors. Not really a similar use, but he had a lot of praise for their devices in general.

• Hamamatsu Prototype detector to the RIKEN design just arrived and ready for tests.

• Intertechnique Still awaiting a device similar to those they built for the MULTICS collaboration which they are loaning us for tests. Had a meeting with Nick Corps from Oxford Instruments, Intertechniques representatives here in the UK. He has gone off to get more information on their capabilities with regard to strip detectors. □